sam (.sam.):

Expand to first 3 terms \[\huge \frac{3+4x+x^2}{\sqrt[3]{1+\frac{x}{2}}}\]

OpenStudy (anonymous):

\[3 + 7/2 +\] to lazy to calculate 3º derivative in 0.

OpenStudy (turingtest):

Taylor series? oh man...

OpenStudy (anonymous):

calculate the second derivative and use taylor series

sam (.sam.):

I don't think so, Use expansion?

sam (.sam.):

\[(1+x)^n=1+nx+\frac{n(n-1)}{2}x^2\]

sam (.sam.):

What i predicted, \[(3+4x+x^2)(1+x/2)^{-1/3}\] then n=-1/3

OpenStudy (anonymous):

u are using binomial formula for not natural exponents....

OpenStudy (anonymous):

i made the taylor expansion in 0

OpenStudy (anonymous):

u can make it anywhere just avoid -2

OpenStudy (turingtest):

I've never expanded a negative fractional exponent befor, I think I'lll have to get back to you... curious to see how to do it though, maybe factoring the top can help somehow

OpenStudy (turingtest):

I tried rationalizing the denominator and simplifying as much as possible, but that was to no avail... still thinking though

sam (.sam.):

ok

sam (.sam.):

@JamesJ

OpenStudy (jamesj):

There's an easier way to do this. Find the Taylor series of \[ \frac{1}{(1+y)^{1/3}} \] Once you have the first three or so terms of that, substitute y = x/2 and multiply that series by (3 + 4x + x^2)

sam (.sam.):

My book didn't teach me about taylor series (even my syllabus does not include taylor series) Is it possible to use \[(1+x)^{n}=1+nx+\frac{n(n-1)}{2}x^2\]?

OpenStudy (jamesj):

Yes

sam (.sam.):

when i use n=-1/3 x=x/2 (bit confusing for x=x/2 part) (3+4x+x^2)(1+x/2)^(−1/3) Expand----------- \[(3+4x+x^2)[1-\frac{x}{6}+\frac{-\frac{1}{3}(-\frac{1}{3}-1)}{2}(\frac{x^2}{4})]\]

sam (.sam.):

My answer is \[\frac{x^4}{18}+\frac{2x^3}{9}-\frac{x^2}{6}+7x\] but thats not the correct answer

OpenStudy (jamesj):

So the terms of 1/(1+(x/2))^(1/3) are 1 - x/6 + x^2/18 as you have it. Double check your multiplication. Notice both have a constant term, but there's no constant in your product of the two.

sam (.sam.):

Oh whops multiplied wrongly, lol so this is the answer \[\frac{x ^{4}}{18}+\frac{x ^{3}}{18}+\frac{x^2}{2}+\frac{7x}{2}+3\] thanks @JamesJ :D